# 1]organization=IIT-CNR, addressline=Via Moruzzi 1, code=56124, city=Pisa, country=Italy

### v3

@@ 8.4 Feature importance
We rely on SHapley Additive exPlanations (SHAP) and Permutation Importance (PI) as the building blocks of our feature importance analysis. PI was introduced in Section 8.2. Instead, SHAP is a method based on cooperative game theory that is widely used for interpreting the output of machine learning models and for estimating the contribution of individual features to a model’s predictions [78, 60]. Leveraging two different post-hoc explainers allows us to gain a deeper understanding of the importance of the implemented features for this task.
[…]
@@ 9.1 Detecting abandoning users
Churn prediction has important social and practical implications [79]. Here we defined and tackled the novel task of predicting churn (i.e., user abandonment) following a massive deplatforming moderation intervention on Reddit [18]. In consideration of the novelty of the task —tackled here for the first time— our results are promising. For example, we achieved micro F1 $=0.914$ when detecting users who immediately halted activity after the moderation intervention (hard abandonment), and micro F1 $=0.814$ when detecting users who initially maintained some activity but who eventually left Reddit (soft abandonment). Albeit preliminary, these results demonstrate that the task of preemptively estimating the effects of future moderation interventions is indeed feasible. Adding to the encouraging findings are the results of our leave-one-out cross-validation (LOOCV) analysis, which underscore the absence of substantial performance degradation when applying our model to unseen data and shows that the model learned generalizable behavioral patterns. At the same time however, our study highlighted some of the challenges of this new task. Among them are the difficulties at forecasting complex and heterogeneous user reactions [19], a dire class imbalance [50], a partial and unreliable view of platform moderation actions [4], and the growing difficulties at accessing platform data [80], which culminate in the lack of extensive labeled datasets. The inherent challenges of predicting the effects of moderation interventions —exemplified by our promising yet preliminary results— imply that much work remains to be done before moderators and platform administrators will be able to leverage powerful and dependable tools for accurately estimating the effects of their actions. To this end, this study paves the way for future endeavors in this emerging area of content moderation, which proposes a new pathway for empowering moderators to optimize interventions based on desired outcomes, such as reducing toxicity while minimizing user attrition and churn. This is currently an as-yet unexplored frontier of moderation, which will open up once it achieves high performance in tasks such as the one proposed and tackled in this work.
@@ 9.2 Characterizing moderated users
Although the performances of the models developed in this study are promising, there is much room for improvement. Model performance strongly depends on the informativeness of the provided features. In this initial study, we computed and tested an extensive and diverse set of 142 features, many of which were borrowed from recent literature on related tasks [81, 31, 63, 60]. Our set of features included information about the activity of the users before the moderation intervention, their participation in multiple communities, their relationships with other users, their use of toxic or otherwise aggressive speech, and their writing style. While some of the extracted features turned out to be powerful predictors in our models, a large share of the features that we evaluated was not particularly informative, which hindered model performance. This suggests that our task, proposed here for the first time, is substantially different from others previously addressed in related literature and for which those features provided robust performance. This finding underscores the importance of tailoring feature selection to the specific characteristics of the task, rather than relying on already-proposed features. Future efforts towards the prediction of moderation intervention effects should thus consider implementing and experimenting with a broader and even more diversified set of features.
@@
Understanding the characteristics that influence user abandonment following a moderation intervention also provides valuable theoretical insights into online community dynamics. Our feature importance analysis highlights key behavioral patterns that distinguish users who leave the platform from those who remain, shedding light on possible psychological and sociological mechanisms underlying their decisions. As discussed in Section 8.4, the trend of comments in non-banned subreddits is among the most informative features. Users who were already decreasing their activity outside the banned communities before the intervention were more likely to abandon the platform entirely, suggesting that their engagement was closely tied to the restricted communities. This aligns with prior research suggesting that individuals with strong ties to specific groups may struggle to integrate into new ones after sudden disruptions [82, 83]. Additionally, the ratio between the number of non-banned and banned subreddits a user participated in proved highly predictive of abandonment. Users who were more deeply embedded in banned communities, with relatively few interactions elsewhere, were more likely to leave. This supports the idea that having alternative engagement opportunities can act as a buffer against platform departure. Indicators of toxicity also emerged as strong predictors of abandonment. Users with higher toxicity levels were more likely to leave after a ban, potentially reflecting a pattern where users who frequently engage in disruptive behaviors feel alienated or unwilling to conform to platform norms following intervention. This observation aligns with theories in digital sociology that suggest enforcement actions can reinforce disengagement among users who strongly identify with rule-violating behaviors [84]. These findings offer preliminary insights into how moderation interventions shape user behavior, with implications for both theoretical understanding and practical applications.
@@
Based on these observations, promising directions for future work involve the development of socio-psychological features. In fact, predicting the effects of moderation interventions essentially revolves around predicting user behavior. Moreover, it is known that socio-psychological characteristics influence online behaviors, including toxic and aggressive ones [19, 85]. A recent demonstration of the latter can be found in those works who measured a significant improvement in hate speech detection tasks, when incorporating socio-psychological features [86]. By the same token, those features could provide valuable information also in the related task of predicting behavioral changes following a moderation intervention, and particularly for those interventions targeting hateful or toxic users.
@@ 9.3 Scalability and Implementation
For real-world deployment, scalability and computational efficiency are key considerations. Our approach leverages traditional machine learning models, which are computationally efficient and significantly less resource-intensive than deep learning-based alternatives. Additionally, many of our features can be precomputed and updated incrementally, reducing the overhead of real-time inference. Notably, our feature importance analysis indicates that only a subset of features is sufficient for accurate predictions, suggesting that models could be further optimized by focusing on the most informative features. Moreover, real-time processing is not a strict requirement, as large-scale moderation interventions (e.g., subreddit bans) are relatively infrequent and do not require continuous, high-speed predictions. Nonetheless, future implementations could explore parallel computing, model compression techniques, or cloud-based deployment to further enhance efficiency.
@@ 9.4 Cross-platform Applicability
The application of our predictive approach to estimating the effects of moderation intervention across different online platforms presents both opportunities and challenges. In principle, our approach is general and its success in a new environment primarily depends on data availability and the ability to leverage platform-informative features. However, each platform has unique conventions, user behaviors, and moderation practices, which affect the generalizability of predictive models. Nonetheless, the existence of platforms featuring similar characteristics provide initial avenues for adaptation. For instance, Reddit shared structural and behavioral similarities with Voat, a now-defunct platform that attracted deplatformed communities. YouTube’s dynamics resemble those of BitChute, where content creators rely heavily on subscriber engagement. Likewise, Twitter and Weibo operate as fast-paced, influencer-driven networks, while TikTok and Instagram emphasize visual content and algorithm-driven exposure. These similarities could suggest that models trained on one platform could be transferred to analogous platforms with relatively minor modifications. At the same time however, major differences between platforms must be carefully considered. Platforms might differ not only in affordances and user interaction patterns, but also in their moderation policies [4]. While Reddit primarily enforces moderation at the community level (e.g., subreddit bans)  [21, 18], Twitter/X and Facebook focus more on individual accounts  [87, 88], and YouTube’s interventions often revolve around demonetization or content takedowns  [89, 90]. These differences imply that a system trained to estimate the effects of a certain type of intervention cannot be directly used to predict the effects of a different intervention, even within the same platform. Each intervention type —whether account suspensions, content removals, or demonetization— affects users in distinct ways, requiring models specifically tailored to the nature of the intervention. Ultimately, the heterogeneous landscape of online platforms and their evolving moderation practices pose challenges to cross-platform applicability. Adapting predictive moderation tools requires both scientific and practical efforts to ensure that models remain accurate and relevant in different digital ecosystems. Future research should explore the trade-off between platform-independent and platform-specific features so as to best capture user engagement and retention dynamics, while also considering how shifting moderation strategies affect the accuracy of predictive models.
@@ 9.5 Predicting effects: classification, quantification, regression
In this work we tackled the problem of predicting the users that would abandon Reddit following a moderation intervention as a binary classification task. Currently, there is a dire lack of tools and systems to predict or estimate the effects of moderation interventions. Within this context, our work paves the way to the development of decision support tools to assist moderators and platform administrators in carrying out effective and efficient content moderation, thus overcoming the limitations of the current trial-and-error approach [13]. In fact, in light of the promising results that we achieved, we envision the possibility to develop additional tools in the near future, to further support moderation endeavors by leveraging established machine learning paradigms. Among them are quantification and regression approaches. For instance, regarding the former, a platform may not be interested in knowing which specific users are likely to abandon or remain on the platform following a moderation intervention. Such a need could arise, among other reasons, due to privacy concerns [5]. In that case, effective moderation could still be guaranteed by computing aggregated estimates of how many users would likely leave compared to those who would remain. From the methodological standpoint, such a problem would be better addressed as a quantification task rather than as a classification one [91]. As a matter of fact, recent results on quantification have demonstrated the superiority of this approach over classify-and-count strategies [92]. Quantification tasks for obtaining aggregated estimates of the effects of moderation interventions thus represent a favorable avenue for future research. Another limitation of classification approaches is the narrow view it offers on post-moderation user behavior. As an example, in order to tackle the binary classification task, we had to come up with a definition for abandoning users. Given the inherent complexity and multifaceted nature of user behavior, we experimented with two different definitions of abandonment —hard and soft, each with its own strengths and weaknesses— so as to account for a broader spectrum of heterogeneous user behavior. Nonetheless, a binary classification task inevitably conceals important details about actual user reactions [14]. For this reason, we envision the possibility to cast the prediction of the effects of moderation interventions as a regression task [63]. This nuanced approach would avoid the need for limiting binary labels in favor of fine-grained estimates of the expected behavioral changes. By moving beyond simplistic labels, such an approach would empower platform moderators with deeper insights into user reactions and patterns, facilitating targeted and effective intervention strategies to foster user retention and community growth. However, solving a regression task with sufficient accuracy poses even more challenges than those addressed here. To this end, additional work on feature engineering is essential to provide the models with more informative features. Moreover, further work should also be directed towards model development, for example by employing sophisticated models capable of fully exploiting the provided features.
@@ 10 Conclusions and Future Work
We proposed and tackled —for the first time— the task of predicting the effects of a moderation intervention. Specifically, we detected those users who abandoned Reddit following the ban of a large number of communities on the platform. To solve this task, we investigated the behavior of 16,540 users by leveraging 16 million comments they posted in the banned communities and 13.8 million comments they posted in non-banned ones. Starting from this extensive dataset, we extracted 142 features conveying information about the activity, toxicity, relations, and writing style of the analyzed users. Our results are promising, albeit preliminary as one would expect from a new task, with the best model achieving micro F1 $=0.914$. Our model proved to be sufficiently generalizable when applied to users from unseen communities. Furthermore, we found that activity features are the most informative for this task, followed by relational and toxicity features. Conversely, writing style features seem to provide limited information.
@@ Acknowledgments
This work is partially supported by the European Union, Next Generation EU, within the PRIN 2022 framework project PIANO (Personalized Interventions Against Online Toxicity); by the PNRR-M4C2 (PE00000013) "FAIR-Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI", funded under the NextGeneration EU program; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).

### v4

@@ 8.4 Feature importance
We rely on SHapley Additive exPlanations (SHAP) and Permutation Importance (PI) as the building blocks of our feature importance analysis. PI was introduced in Section 8.2. Instead, SHAP is a method based on cooperative game theory that is widely used for interpreting the output of machine learning models and for estimating the contribution of individual features to a model’s predictions [lundberg2017unified, mazza2022investigating]. Leveraging two different post-hoc explainers allows us to gain a deeper understanding of the importance of the implemented features for this task.
To obtain the overall importance of each feature for our task, we use an information fusion-based sensitivity analysis technique [tardelli2022detecting]. Concretely, SHAP estimates feature importance at the individual data point level, requiring values to be aggregated across the entire dataset. Therefore, for each feature $f_{i}$ and for each trained model $M_{j}\in\{\texttt{NB},\texttt{KNN},\texttt{DT},\texttt{RF},\texttt{AB},\texttt{GB},\texttt{SVM}\}$, we use SHAP to compute the contribution $C(i,j,d)$ of $f_{i}$ to the prediction given by $M_{j}$ for a specific data point $d$. We then aggregate the contributions over all data points $N$ in the test set to compute a local score:
@@ 9.1 Detecting abandoning users
Churn prediction has important social and practical implications [panimalar2023customer]. Here we defined and tackled the novel task of predicting churn (i.e., user abandonment) following a massive deplatforming moderation intervention on Reddit [cima2025investigating]. In consideration of the novelty of the task —tackled here for the first time— our results are promising. For example, we achieved micro F1 =0.914=0.914 when detecting users who immediately halted activity after the moderation intervention (hard abandonment), and micro F1 =0.814=0.814 when detecting users who initially maintained some activity but who eventually left Reddit (soft abandonment). Albeit preliminary, these results demonstrate that the task of preemptively estimating the effects of future moderation interventions is indeed feasible. Adding to the encouraging findings are the results of our leave-one-out cross-validation (LOOCV) analysis, which underscore the absence of substantial performance degradation when applying our model to unseen data and shows that the model learned generalizable behavioral patterns. At the same time however, our study highlighted some of the challenges of this new task. Among them are the difficulties at forecasting complex and heterogeneous user reactions [cresci2022personalized], a dire class imbalance [robertson2022uncommon], a partial and unreliable view of platform moderation actions [trujillo2023dsa], and the growing difficulties at accessing platform data [jaursch2024dsa], which culminate in the lack of extensive labeled datasets. The inherent challenges of predicting the effects of moderation interventions —exemplified by our promising yet preliminary results— imply that much work remains to be done before moderators and platform administrators will be able to leverage powerful and dependable tools for accurately estimating the effects of their actions. To this end, this study paves the way for future endeavors in this emerging area of content moderation, which proposes a new pathway for empowering moderators to optimize interventions based on desired outcomes, such as reducing toxicity while minimizing user attrition and churn. This is currently an as-yet unexplored frontier of moderation, which will open up once it achieves high performance in tasks such as the one proposed and tackled in this work.
@@ 9.2 Characterizing moderated users
Although the performances of the models developed in this study are promising, there is much room for improvement. Model performance strongly depends on the informativeness of the provided features. In this initial study, we computed and tested an extensive and diverse set of 142 features, many of which were borrowed from recent literature on related tasks [varol2018feature, habib2022proactive, tardelli2022detecting, mazza2022investigating]. Our set of features included information about the activity of the users before the moderation intervention, their participation in multiple communities, their relationships with other users, their use of toxic or otherwise aggressive speech, and their writing style. While some of the extracted features turned out to be powerful predictors in our models, a large share of the features that we evaluated was not particularly informative, which hindered model performance. This suggests that our task, proposed here for the first time, is substantially different from others previously addressed in related literature and for which those features provided robust performance. This finding underscores the importance of tailoring feature selection to the specific characteristics of the task, rather than relying on already-proposed features. Future efforts towards the prediction of moderation intervention effects should thus consider implementing and experimenting with a broader and even more diversified set of features.
@@
Understanding the characteristics that influence user abandonment following a moderation intervention also provides valuable theoretical insights into online community dynamics. Our feature importance analysis highlights key behavioral patterns that distinguish users who leave the platform from those who remain, shedding light on possible psychological and sociological mechanisms underlying their decisions. As discussed in Section 8.4, the trend of comments in non-banned subreddits is among the most informative features. Users who were already decreasing their activity outside the banned communities before the intervention were more likely to abandon the platform entirely, suggesting that their engagement was closely tied to the restricted communities. This aligns with prior research suggesting that individuals with strong ties to specific groups may struggle to integrate into new ones after sudden disruptions [brown2000social, hu2024social]. Additionally, the ratio between the number of non-banned and banned subreddits a user participated in proved highly predictive of abandonment. Users who were more deeply embedded in banned communities, with relatively few interactions elsewhere, were more likely to leave. This supports the idea that having alternative engagement opportunities can act as a buffer against platform departure. Indicators of toxicity also emerged as strong predictors of abandonment. Users with higher toxicity levels were more likely to leave after a ban, potentially reflecting a pattern where users who frequently engage in disruptive behaviors feel alienated or unwilling to conform to platform norms following intervention. This observation aligns with theories in digital sociology that suggest enforcement actions can reinforce disengagement among users who strongly identify with rule-violating behaviors [thomas2021behavior]. These findings offer preliminary insights into how moderation interventions shape user behavior, with implications for both theoretical understanding and practical applications.
@@
Based on these observations, promising directions for future work involve the development of socio-psychological features. In fact, predicting the effects of moderation interventions essentially revolves around predicting user behavior. Moreover, it is known that socio-psychological characteristics influence online behaviors, including toxic and aggressive ones [cresci2022personalized, giorgi2025human]. A recent demonstration of the latter can be found in those works who measured a significant improvement in hate speech detection tasks, when incorporating socio-psychological features [raut2023enhancing]. By the same token, those features could provide valuable information also in the related task of predicting behavioral changes following a moderation intervention, and particularly for those interventions targeting hateful or toxic users.
training time (s) test time (s)
task model p F1 total features avg total
hard abandonment KNN 20 0.384 0.003 1,138.809 0.003 9.553
NB 10 0.297 0.062 1,131.399 0.001 4.696
RF 142 0.675 45.976 2,540.524 0.018 57.311
AB 142 0.659 10.648 2,540.524 0.011 37.132
DT 10 0.430 0.062 1,132.099 0.001 3.982
GB 142 0.695 36.877 2,540.524 0.005 16.483
SVM 50 0.434 6.443 2,521.658 0.003 9.195
soft abandonment KNN 10 0.473 0.021 1,101.728 0.002 7.5051
NB 10 0.438 0.007 1,101.728 0.002 4.948
RF 142 0.578 11.612 2,540.524 0.007 24.189
AB 142 0.588 11.693 2,540.524 0.012 40.316
DT 20 0.494 0.234 1,136.813 0.002 5.469
GB 142 0.590 23.668 2,540.524 0.005 16.270
SVM 50 0.499 11.189 1,155.545 0.003 10.460
Table 7: Training and test times (in seconds) for each model across both abandonment tasks. Test performance is reported in terms of time to extract the features (features), average prediction time per user (avg), and total prediction time (total). Since training is performed once on the full dataset, only the total training time (total) is reported. Additionally, we report the number of features used (p) and the resulting F1 score of each model.
@@ 9.3 Scalability and Implementation
For real-world deployment, scalability and computational efficiency are key considerations. Our approach leverages traditional machine learning models, which are computationally efficient and significantly less resource-intensive than deep learning-based alternatives. Additionally, many of our features can be precomputed and updated incrementally, reducing the overhead of real-time inference. Notably, our feature importance analysis indicates that only a subset of features is sufficient for accurate predictions, suggesting that models could be further optimized by focusing on the most informative features.
To estimate the computational cost of our approach in a real-world scenario, we trained and tested our machine learning models on a training set of 12,994 users, a test set of 3,249 users, and a total of 1.57M comments. We measured both training and inference times for the two abandonment tasks. All experiments were conducted on a virtual machine running Ubuntu 20.04 with 32 GB of RAM and 16 virtual CPUs (Intel Xeon, 2.2 GHz). The results are reported in Table 7. As expected, across both tasks, simpler models exhibit the lowest training and test times, whereas the ensemble models require significantly more, up to several orders of magnitude in the worst cases. This is attributable to their greater complexity and the considerably higher number of features they use. Nonetheless, there is a significant trade-off in terms of F1 score, given that Gradient Boosting (GB), AdaBoost (AB), and Random Forest (RF) emerge as the best-performing models. The feature extraction time is significantly high for all models, including those using the lowest number of features. This is due to the reliance on a subset of computationally expensive features, such as part-of-speech tagging and toxicity scores, which dominate the overall extraction cost. However, in our specific case real-time processing is not a strict requirement as large-scale moderation interventions (e.g., subreddit bans) are relatively infrequent and do not require continuous, high-speed predictions. Therefore, it is feasible to favor more complex models in exchange for better predictive performance. Results reported in Table 7 are also valuable to highlight the trade-off between model efficiency and effectiveness. To conclude, future implementations could explore parallel computing, model compression techniques, or cloud-based deployment to further enhance efficiency.
@@ 9.4 Cross-platform Applicability
The application of our predictive approach to estimating the effects of moderation intervention across different online platforms presents both opportunities and challenges. In principle, our approach is general and its success in a new environment primarily depends on data availability and the ability to leverage platform-informative features. However, each platform has unique conventions, user behaviors, and moderation practices, which affect the generalizability of predictive models. Nonetheless, the existence of platforms featuring similar characteristics provide initial avenues for adaptation. For instance, Reddit shared structural and behavioral similarities with Voat, a now-defunct platform that attracted deplatformed communities. YouTube’s dynamics resemble those of BitChute, where content creators rely heavily on subscriber engagement. Likewise, Twitter and Weibo operate as fast-paced, influencer-driven networks, while TikTok and Instagram emphasize visual content and algorithm-driven exposure. These similarities could suggest that models trained on one platform could be transferred to analogous platforms with relatively minor modifications. At the same time however, major differences between platforms must be carefully considered. Platforms might differ not only in affordances and user interaction patterns, but also in their moderation policies [trujillo2023dsa]. While Reddit primarily enforces moderation at the community level (e.g., subreddit bans)  [trujillo2021echo, cima2025investigating], Twitter/X and Facebook focus more on individual accounts  [pierri2023does, kalsnes2021hiding], and YouTube’s interventions often revolve around demonetization or content takedowns  [zappin2022youtube, ma2022m]. These differences imply that a system trained to estimate the effects of a certain type of intervention cannot be directly used to predict the effects of a different intervention, even within the same platform. Each intervention type —whether account suspensions, content removals, or demonetization— affects users in distinct ways, requiring models specifically tailored to the nature of the intervention. Ultimately, the heterogeneous landscape of online platforms and their evolving moderation practices pose challenges to cross-platform applicability. Adapting predictive moderation tools requires both scientific and practical efforts to ensure that models remain accurate and relevant in different digital ecosystems. Future research should explore the trade-off between platform-independent and platform-specific features so as to best capture user engagement and retention dynamics, while also considering how shifting moderation strategies affect the accuracy of predictive models.
@@ 9.5 Predicting effects: classification, quantification, regression
In this work we tackled the problem of predicting the users that would abandon Reddit following a moderation intervention as a binary classification task. Currently, there is a dire lack of tools and systems to predict or estimate the effects of moderation interventions. Within this context, our work paves the way to the development of decision support tools to assist moderators and platform administrators in carrying out effective and efficient content moderation, thus overcoming the limitations of the current trial-and-error approach [trujillo2022make]. In fact, in light of the promising results that we achieved, we envision the possibility to develop additional tools in the near future, to further support moderation endeavors by leveraging established machine learning paradigms. Among them are quantification and regression approaches. For instance, regarding the former, a platform may not be interested in knowing which specific users are likely to abandon or remain on the platform following a moderation intervention. Such a need could arise, among other reasons, due to privacy concerns [gorwa2020algorithmic]. In that case, effective moderation could still be guaranteed by computing aggregated estimates of how many users would likely leave compared to those who would remain. From the methodological standpoint, such a problem would be better addressed as a quantification task rather than as a classification one [gonzalez2017review]. As a matter of fact, recent results on quantification have demonstrated the superiority of this approach over classify-and-count strategies [esuli2023learning]. Quantification tasks for obtaining aggregated estimates of the effects of moderation interventions thus represent a favorable avenue for future research. Another limitation of classification approaches is the narrow view it offers on post-moderation user behavior. As an example, in order to tackle the binary classification task, we had to come up with a definition for abandoning users. Given the inherent complexity and multifaceted nature of user behavior, we experimented with two different definitions of abandonment —hard and soft, each with its own strengths and weaknesses— so as to account for a broader spectrum of heterogeneous user behavior. Nonetheless, a binary classification task inevitably conceals important details about actual user reactions [trujillo2023one]. For this reason, we envision the possibility to cast the prediction of the effects of moderation interventions as a regression task [tardelli2022detecting]. This nuanced approach would avoid the need for limiting binary labels in favor of fine-grained estimates of the expected behavioral changes. By moving beyond simplistic labels, such an approach would empower platform moderators with deeper insights into user reactions and patterns, facilitating targeted and effective intervention strategies to foster user retention and community growth. However, solving a regression task with sufficient accuracy poses even more challenges than those addressed here. To this end, additional work on feature engineering is essential to provide the models with more informative features. Moreover, further work should also be directed towards model development, for example by employing sophisticated models capable of fully exploiting the provided features.
@@ 10 Conclusions and Future Work
We proposed and tackled —for the first time— the task of predicting the effects of a moderation intervention. Specifically, we detected those users who abandoned Reddit following the ban of a large number of communities on the platform. To solve this task, we investigated the behavior of 16,540 users by leveraging 16 million comments they posted in the banned communities and 13.8 million comments they posted in non-banned ones. Starting from this extensive dataset, we extracted 142 features conveying information about the activity, toxicity, relations, and writing style of the analyzed users. Our results are promising, albeit preliminary as one would expect from a new task, with the best model achieving micro F1 =0.914=0.914. Our model proved to be sufficiently generalizable when applied to users from unseen communities. Furthermore, we found that activity features are the most informative for this task, followed by relational and toxicity features. Conversely, writing style features seem to provide limited information.
@@ Acknowledgments
This work is partially supported by the European Union – Next Generation EU, Mission 4 Component 1, for project PIANO (CUP B53D23013290006); by the PNRR-M4C2 (PE00000013) "FAIR-Future Artificial Intelligence Research" - Spoke 1 "Human-centered AI", funded under the NextGeneration EU program; and by the Italian Ministry of Education and Research (MUR) in the framework of the FoReLab project (Departments of Excellence).